If the time for reaching a steady value is very long, what effect might this have on the measured stopping voltage values? (Photoelectric Effect Lab)
Answer #1

The photoelectric current in a typical setup is extremely small, and making a precise measurement is difficult. Normally the electrons will reach the anode of the photo diode, and their number can be measured from the (minute) anode current. However, we can apply a reverse voltage to the anode; this reverse voltage repels the electrons and prevents them from reaching the anode. The minimum required voltage is called the stopping potential Vs, and the “stopping energy” of each electron is therefore e*Vs. Thus,

e*Vs=hf−W

or

Vs=(h/e)f−W/e

when the steady time is very long it means the stopping potential is very so overcoming the higher stopping potential it takes long time.